OpenStudy (anonymous):

On your birthday, you deposit $540.00 in an account that pays 6% interest compound anually. How much is in the account 3 years later?

OpenStudy (johnweldon1993):

Simple Interest Formula \[\large P = P_0(1 + r)^t\] Where Po is your original amount (540 here) r is your rate...turn your percent into a decimal here (.06 here) and 't' is the time in years (3 here)

OpenStudy (anonymous):

So i multiply it?

OpenStudy (anonymous):

Here is the options : $637.00 $543.18 $643.00 $1,717.20 And if I was suppose to multiply it the answer is not up there.

OpenStudy (johnweldon1993):

So we have \[\large P = 540(1.06)^3\] \[\large P = 540(1.191016)\] \[\large P = ?\]

OpenStudy (anonymous):

Am I supposed to multiply the second one?

OpenStudy (johnweldon1993):

Well yes... \[\large 540(1.191016)\] is the same as \[\large 540 \times 1.191016\]

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

is it c 643.00

OpenStudy (johnweldon1993):

That is correct

OpenStudy (anonymous):

Thanks